40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  SUS Lab eLog  Not logged in ELOG logo
Message ID: 386     Entry time: Thu Jan 26 15:48:16 2012
Author: Giordon Stark 
Type: DailyProgress 
Category: Coating Q 
Subject: Scrounging of Parts and slight issues 

 Zach and Giordon went to the 40m yesterday afternoon and rounded up a whole slew of parts. These include:

  • 3 metallic/silver mirrors
  • 1 633nm Polarizing Beamsplitter (verified wavelength using a pen laser) [PBS]
  • 1 PLCX-25.4-33.7-UV-1064 (planar convex mirror)
  • 1 PLCX-25.4-46.4-UV (planar convex mirror)
  • 1 633nm L/4 (waveplate: quarter-wavelength) [L/4]
  • 1 633nm L/2 (waveplate: half-wavelength) [L/2]

8 items/parts total.

We went back over to the lab in Bridge and took out the parts. There is a laser set-up in there at around 633nm and so we did the following:

  1. Set up the wave plates (L/4 then L/2) and then the PBS
  2. Changed the rotation on L/4 and L/2 and saw that we changed the laser output intensity (it varied) but it never zeroed out
  3. Removed the wave plates, put in L/4 only
  4. Repeated step 2 - saw the output never zeroed out
  5. Removed L/4, put in L/2 only
  6. Repeated step 2 - saw the output never zeroed out

Our conclusion of what's wrong is unknown. Some possible ideas include the laser is not "fixed" (it changes polarizations, it's continuously rotating, etc...). Further testing will be done Friday morning to determine whether the wave plates are not working the way we expect them to or if the laser is the issue.

An updated version of the Experimental Set up with more details will be added after Friday's conclusion.

ELOG V3.1.3-